OpenStudy (anonymous):

Solve for x:

OpenStudy (anonymous):

x:\[\log_{3}(x^2-4)-\log_{3}(x+2)=2 \]

Nnesha (nnesha):

remember \[\log a - \log b = \log \frac{ a }{ b }\] and \[\log a + \log b = \log (a \times b)\]

Nnesha (nnesha):

so how would you change \[\log _{3} (x^2 -4 ) - \log _{3} (x+2) = ????\]

jagr2713 (jagr2713):

First we combine logarithms

jagr2713 (jagr2713):

(log(x^2-4))/(log(3))-(log(x+2))/(log(3)) = (log(1/(x+2)))/(log(3))+(log(x^2-4))/(log(3)) = (log(1/(x+2))+log(x^2-4))/(log(3)) = (log((x^2-4)/(x+2)))/(log(3)): (log((x^2-4)/(x+2)))/(log(3)) = 2

jagr2713 (jagr2713):

Then we multiply both sides by a constant to simplify the equation. Multiply both sides by log(3): log((x^2-4)/(x+2)) = 2 log(3)

jagr2713 (jagr2713):

Then we simplify logarithmic terms on the right hand side. 2 log(3) = log(3^2) = log(9): log((x^2-4)/(x+2)) = log(9)

OpenStudy (anonymous):